EFOSC News Updates

On this page, we list the main changes that affected the instrument and this web site.

      

  • Absorption features in EFOSC2 internal flat spectra. See report (28/07/2013)
  • Filter Gunn r#786 decommisioned:Replaced by Gunn r#784. Gunn r#787 is missing. (16/10/2009)
  • 2nd order effect: We have tested a number of order sorting (high-pass) filters for cutting out the 2nd order effect seen in spectra of blue sources with some grisms. See the new order sorting filters page for details. (26/08/2009)
  • Calibration checks: A new What is normal? page has been created for users to perform a quick visual check of calibration frames. (26/08/2009)
  • Defocused photometry test: We tested the suitability of EFOSC and the NTT to perform high precision relative photometry by extreme telescope defocusing, as has been succesfully done at the Danish 1.54m telescope. See this report. (26/08/2009)
  • Standard Star OBs: New OBs are available for import into p2pp for observation of Landolt photometric standard star fields. They are described in this document. (26/01/2009)
  • New ZP tool: A new script to automatically measure zeropoints following observation of standard stars is available. It is described, along with updated values for the zeropoints, on the EFOSC2 zeropoint page. (23/11/2008)
  • Web pages update: These web pages have been completely revised following the move to the NTT and the move of the La Silla web site to the new ESO "look and feel". (04/11/2008)
  • New manual: An updated version of the manual with information on EFOSC2 at the NTT is ready for P82, following the successful recommissioning of the instrument at the NTT Nasmyth B focus. (21/06/2008)
  • Move to the NTT: EFOSC2 was successfully moved to its new home at the NTT Nasmyth B focus, where it is permanently mounted (April 2008).
  • VPHG grisms: Two new Volume Phase Holographic Grisms now give EFOSC2 a higher resolution spectroscopy mode. See report and Saviane & Monaco. 2008, ESO Messenger 132, 20 (March 2008).
  • Grism #10 decommissioned: Grism #10 (6280-8200 Å) suffered a major damage and is no longer offered. (14/06/2007)
  • Pattern in Bias Frames: A pattern of probable electronic nature is sporadically observed in Bias frames taken with binning 2x2. For details, see Pattern in Bias 2x2. (22/02/2007)
  • Quarter-wave plate: A new polarization unit will be offered in P79; it will allow to measure circular polarimetry, thanks to a quarter-lambda retarder plate (14/08/2006)
  • Calibration plan: EFOSC2 calibration plan now on-line (10/05/2006)
  • Quick-look tool: A short manual of the quick-look tool for long-slit spectroscopy is now available (27/04/2006)
  • New Grisms: All broken grisms have been either repaired or replaced, the last being grism #15 replaced by #17. The new efficiencies and further info can be found at the EFOSC grisms page. (19/11/2004)
  • Fringing pattern determined: The fringing patterns for the filter i#705 has been determined and put on this page. (17/11/2004)
  • Zero-Point monitoring page updated: The zero-points are measured in each photometric night when EFOSC is mounted. The results are stored in the EFOSC ZP-page. (13/11/2004)
  • New Grism Efficiencies: System efficiencies have been measured for all available grisms. For details, see EFOSC grims page. (31/08/2004)